### v2

# Large-order perturbation theory of linear eigenvalue problems

Abstract
We consider a class of linear eigenvalue problems depending on a small parameter $\epsilon$ in which the series expansion for the eigenvalue in powers of $\epsilon$ is divergent. We develop a new technique to determine the precise nature of this divergence. We illustrate the technique through its application to four examples: the anharmonic oscillator, a simplified model of equatorially-trapped Rossby waves, and two simplified models based on quasinormal modes of Reissner-Nordström-de Sitter black holes.
Keywords
Exponential asymptotics; Stokes phenomenon; divergent expansion; WKB theory; quantum mechanics.
1 Introduction
Linear eigenvalue problems of the form
$L_{\epsilon}g=\lambda g,$
where $L_{\epsilon}$ is a differential operator (depending on a parameter $\epsilon>0$), are ubiquitous in applied mathematics and theoretical physics. The eigenvalue $\lambda$ might correspond to an energy level, the frequency of a normal mode of oscillation, or the growth rate in a linear stability analysis. Often $\epsilon$ is a small parameter, in which case it is common to develop the perturbation series for the eigenvalue $\lambda$ in powers of $\epsilon$,
$\lambda\sim\sum_{n=0}^{\infty}\epsilon^{n}\lambda_{n}.$
Sometimes this series diverges, and it is of interest to determine the nature of this divergence. From a purely numerical point of view understanding the large-order behaviour tells you how many terms to take in an optimal approximation, how large the smallest achievable error is, and whether the series can be meaningfully resummed. However, as has long been known, large-order growth is controlled by non-perturbative effects that are not visible in ordinary perturbation theory. For eigenvalue problems, this can mean an exponentially small imaginary part of the eigenvalue, indicating an instability not visible in the regular perturbation expansion, or, in quantum mechanics, tunnelling between classically distinct regions, or exponentially-small level splittings. Here we provide a straightforward approach to determining the precise asymptotic behaviour of $\lambda_{n}$ for large $n$.
One of the earliest examples of such a problem is the quantum anharmonic oscillator [2, 3], which we revisit in Section 3. This problem has a long history; it was the first non-exactly-solvable problem tackled by the newly-written Schrodinger equation in 1926, has practical applications ranging from quantum chemistry and atomic-molecular physics to crystal lattice vibrations in solid-state theory, and serves as a simple model for quantum field theory [13]. The seminal work by Bender and Wu [2, 3] established nature of the divergence of the perturbation series for the ground state energy. This work became the prototype for similar analyses in many other quantum mechanical systems, in what is now known as large-order perturbation theory [1, 10].
The main technique of Bender and Wu is to analytically continue in $\epsilon$ (typically until $\epsilon$ is negative), and then solve the resulting problem by combination of Liouville-Green (WKB) and matched asymptotic approximations. This is a delicate procedure, since the goal is to identify an exponentially small component of $\lambda$ beyond-all-orders of the divergent asymptotic series. Cauchy’s integral formula is then used to determine the coefficients in the power series expansion of $\lambda$ on the positive real $\epsilon$ axis in terms of the values of $\lambda$ on either side of the negative real $\epsilon$ axis. This technique is ingenious, and has proved successful, but the details can be very complicated. The present work aims to present an alternative method, which we hope will be useful.
Intriguing also is the work of Dunne & Ünsal [8, 9], who use a uniform WKB approach for quantum mechanical systems with degenerate minima, focusing in particular on double-well and sine-Gordon potentials. They write the wave function in terms of a scaled parabolic cylinder function $\Psi(x)=D_{\nu}(u(x)/\sqrt{\epsilon})/\sqrt{u^{\prime}(x)}$, where the index $\nu$ is a parameter to be determined along with $u(x)$. Although $u(x)$ and $\lambda$ can be determined locally in terms of $\nu$ by a regular perturbation expansion, $\nu$ itself is determined by imposing a global boundary condition (a symmetry condition at the midpoint between wells). This shows $\nu$ is exponentially close to an integer, and provides an exponential correction to $\lambda$. In [8] this exponential correction is used to determine the late-order terms in the power series, in a similar way to Bender & Wu.
In the applied mathematics literature an early example of such a problem occurs in the work of Boyd and Natarov [4], who consider a model problem for an equatorially-trapped Rossby wave in a shear flow in the ocean or atmosphere. There the main interest is in the imaginary part of the eigenvalue (corresponding to the growth rate of instability)—the divergent perturbation series is purely real, but there is an exponentially small imaginary part beyond all orders. In [12] this problem is attacked in almost the reverse direction to Bender and Wu—the divergent series is first found, and used to determine the exponentially small imaginary component of the eigenvalue via optimal truncation and Stokes phenomenon, rather than the other way round. Simplifying and extending the procedure from [12] forms the basis of the present work.
We present our procedure through its application to four examples. Each follows the same general framework, but the final part of the analysis differs slightly in each case. We hope this will allow the interested reader to adapt the method to their own particular problem.
The general framework is:
(i) Inner region. Each problem starts with a regular perturbation expansion. Typically the coefficients in the expansion of the eigenvalue are determined by imposing a regularity condition at the origin $x=0$, and each term in the expansion is a polynomial in $x$. This region determines the eigenvalue expansion completely, but it is difficult to extract the large-order behaviour from the resulting recurrence relations.
(ii) Outer region. The regular expansion in (i) is not uniform, and rearranges for large $x$, leading to a new expansion once $x=X/\epsilon$ has been rescaled. The outer problem is a singular perturbation problem, so that this expansion diverges in the usual form of factorial/power, driven by singularities away from $X=0$. The large-order behaviour of this divergence is easy to determine by now standard methods. In addition the late terms have an independent component driven by the divergent eigenvalue expansion.
(iii) Boundary layer in the late terms near $X=0$. This is the new ingredient to the method. The key observation is that the large-order approximation of the outer expansion (ii) is also non-uniform, so that there is another inner region near $X=0$, now not in the small-$\epsilon$ expansion but in the large-order expansion. The resolution of this inner region links the two parts of the expansion in (ii), and determines the large-order behaviour of the eigenvalue.
2 Example 1: Simplified black holes
We consider the model problem
$2(1-\epsilon x)(-\omega g+xg^{\prime})+g+(xg^{\prime})^{\prime}=0,\qquad-\infty<x<0,$ (1)
with $g(0)=1$ and $g(x)=o(\mathrm{e}^{-x})$ as $x\rightarrow$-$\infty$, where $0<\epsilon\ll 1$. This is a much-simplified version of the problem in [7] concerning quasinormal modes of Reissner-Nordström-de Sitter black holes, keeping only those ingredients necessary to illustrate the methodology; very roughly speaking $g$ is the charged scalar field perturbation, $x$ is the radial distance measured from the cosmological horizon, $\epsilon$ is the charge, and the eigenvalue $\omega$ is the frequency of the mode.
As $x\rightarrow$-$\infty$ the two possible behaviours are
$g(x)\sim\mathrm{e}^{\epsilon x^{2}},\qquad g(x)\sim x^{\omega},$
while as $x\rightarrow 0$ the two possible behaviours are
$g(x)\sim 1,\qquad g(x)\sim\log x.$
The boundary conditions at $x=0$ and $x=$-$\infty$ each remove one degree of freedom, so that there is a nonzero solution only if $\omega$ takes particular values. The goal is to find the asymptotic expansion of the eigenvalues,
$\omega\sim\sum_{n=0}^{\infty}\epsilon^{n}\omega_{n},$
as $\epsilon\rightarrow 0$, and in particular the form of the divergence of $\omega_{n}$ as $n\rightarrow\infty$. To simplify the presentation we will focus on the leading eigenvalue, here and in each of our examples, but of course the methodology works for any eigenvalue.
2.1 Inner region
We start with
$2(1-\epsilon x)(-\omega g+xg^{\prime})+g+(xg^{\prime})^{\prime}=0.$
We expand
$g=\sum_{n=0}^{\infty}\epsilon^{n}g_{n},\qquad\omega=\sum_{n=0}^{\infty}\epsilon^{n}\omega_{n},$ (2)
to give at leading order
$2(-\omega_{0}g_{0}+xg_{0}^{\prime})+g_{0}+(xg_{0}^{\prime})^{\prime}=0.$
The solution which is regular at the origin is
$g_{0}=L_{\omega_{0}-1/2}(-2x),$
where $L_{n}(z)$ is the Laguerre function. To avoid exponential growth as $x\rightarrow$-$\infty$ we need the Laguerre function to be a polynomial, i.e. we need $\omega_{0}-1/2$ to be a non-negative integer. Choosing the first of these, $n=0$, gives the solution $g_{0}=1$, $\omega_{0}=1/2$. At next order
$2xg_{1}^{\prime}+(xg_{1}^{\prime})^{\prime}-2\omega_{1}=-x.$
The solution which is regular at $x=0$ and does not grow exponentially at minus infinity is
$g_{1}=-\frac{x}{2},\qquad\omega_{1}=-\frac{1}{4}.$
In general
$2xg_{n}^{\prime}+(xg_{n}^{\prime})^{\prime}-2\omega_{n}=-2x(\omega_{n-1}g_{0}+\cdots+\omega_{0}g_{n-1}-xg_{n-1}^{\prime})+2(\omega_{n-1}g_{1}+\cdots\omega_{1}g_{n-1}),$ (3)
and the solution is of the form
$g_{n}=\sum_{i=1}^{n}a_{ni}x^{i},$
with
$2ja_{n,j}+(j+1)^{2}a_{n,j+1}$ $=$ $-2\sum_{k=j-1}^{n-1}\omega_{n-1-k}a_{k,j-1}+2(j-1)a_{n-1,j-1}+2\sum_{k=j}^{n-1}\omega_{n-k}a_{k,j},$ (4)
 $\mbox{ }\hskip 113.81102pt\qquad\mbox{ for }j=1,\ldots,n,$
$a_{n1}-2\omega_{n}$ $=$ $0.$ (5)
We can iterate to find $\omega_{n}$ numerically. Figure 1(a) shows $|\omega_{n}|^{1/n}$ as a function of $n$; the linear growth in $n$ is consistent with factorial growth in $\omega_{n}$ at large $n$. In principle we could extract the asymptotic behaviour as $n\rightarrow\infty$ from (4)-(5), but this is not so straightforward. The method we now highlight determines $\omega_{n}$ for large $n$ without the need to analyse (4)-(5).
2.2 Outer region
The expansion (2) is not uniform in xx—it rearranges when $x$ is large. In this section we develop the corresponding expansion valid for large $x$.
To this end we set $\epsilon x=X$ to give
$2(1-X)(-\omega g+Xg^{\prime})+g+\epsilon(Xg^{\prime})^{\prime}=0.$
Now expanding
$g=\sum_{n=0}^{\infty}\epsilon^{n}g_{n}$ (6)
gives, at leading order,
$2(1-X)(-\omega_{0}g_{0}+Xg_{0}^{\prime})+g_{0}=0,$
so that
$g_{0}=B(1-X)^{1/2}X^{\omega_{0}-1/2},$
for some constant $B$. For there to be no singularity at $X=0$ we require $\omega_{0}=1/2$, in agreement with §2.1. To match with the inner expansion as $X\rightarrow 0$ we require $g_{0}\rightarrow 1$ so that $B=1$. In general, equating coefficients of $\epsilon^{n}$,
$X(g_{n}+2(1-X)g_{n}^{\prime})=-(Xg_{n-1}^{\prime})^{\prime}+2(1-X)(\omega_{1}g_{n-1}+\cdots+\omega_{n}g_{0}).$ (7)
We need to determine the late terms in the expansion, that is, the behaviour of $g_{n}$ as $n\rightarrow\infty$. There are two sources of divergence in $g_{n}$: the usual factorial/power divergence driven by differentiating $g_{n-1}$, and a factorial/constant divergence driven by $\omega_{n}$. For the first, we follow the usual procedure [5] by supposing that
$g_{n}\sim\frac{G\Gamma(n+\gamma)}{\chi^{n+\gamma}}$ (8)
as $n\rightarrow\infty$, where $G$ and $\chi$ are functions of $x$ and $\gamma$ is constant. Then, equating coefficients of powers of $n$ gives, at leading order,
$\chi^{\prime}=2(1-X).$
Since this divergence is driven by the singularity in $g_{0}$ at $X=1$, we have $\chi(1)=0$, so that
$\chi=-(1-X)^{2}.$
At next order we find
$(2-5X)G+2X(1-X)G^{\prime}=0,$
giving
$G=\frac{\Lambda}{X(1-X)^{3/2}},$
for some constant $\Lambda$. Thus (absorbing $(-1)^{-\gamma}$ into $\Lambda$) this part of $g_{n}$ satisfies
$g_{n}\sim\frac{\Lambda(-1)^{n}\Gamma(n+\gamma)}{X(1-X)^{3/2}(1-X)^{2n+2\gamma}}.$
As $X\rightarrow 1$,
$g_{n}\sim\frac{\Lambda(-1)^{n}\Gamma(n+\gamma)}{(1-X)^{3/2}(1-X)^{2n+2\gamma}}.$
Comparing powers of 1X1-X with $g_{0}$ gives
$\frac{3}{2}-2\gamma=\frac{1}{2}\qquad\Rightarrow\qquad\gamma=-1,$
so that
$g_{n}\sim\frac{\Lambda(-1)^{n}\Gamma(n-1)}{X(1-X)^{2n-1/2}}.$ (9)
The other part of $g_{n}$, driven by the divergence of $\omega_{n}$, is given by $g_{n}\sim Q\omega_{n}$ where
$X(Q+2(1-X)Q^{\prime})=2(1-X)g_{0}=2(1-X)^{3/2},$
so that
$Q=(1-X)^{1/2}(\log X+C).$
The presence of $\log X$ here means we need to modify slightly the ansatz $g_{n}\sim Q\omega_{n}$ (essentially we need $C$ to include a term proportional to $\log n$). If we set instead $g_{n}=(Q_{0}\log n+Q_{1})\omega_{n}$ then
$X(Q_{0}-2(1-X)Q_{0}^{\prime})$ $=$ $0,$
$X(Q_{1}-2(1-X)Q_{1}^{\prime})$ $=$ $2(1-X)^{3/2},$
so that
$Q_{0}\log n+Q_{1}=(1-X)^{1/2}\left(C_{0}\log n+C_{1}+\log X\right).$
Putting the two parts of $g_{n}$ together gives
$g_{n}\sim\frac{\Lambda(-1)^{n}\Gamma(n-1)}{X(1-X)^{2n-1/2}}+(1-X)^{1/2}(C_{0}\log n+C_{1}+\log X)\omega_{n}.$ (10)
To determine $\Lambda$ we need to match with an inner region in the vicinity of the singularity at $X=1$.
2.3 Inner region near $X=1$
Motivated by both $g_{0}(X)=\sqrt{1-X}$ and by (9) we set $X=1-\epsilon^{1/2}\hat{x}$, $g=\epsilon^{1/4}\hat{g}$ to give
$(1-2\hat{x}\epsilon^{1/2}\omega)\hat{g}+(2\hat{x}(-1+\epsilon^{1/2}\hat{x})-\epsilon^{1/2})\hat{g}^{\prime}+(1-\epsilon^{1/2}\hat{x})\hat{g}^{\prime\prime}=0.$ (11)
In terms of the inner variable
$g_{0}$ $=$ $\epsilon^{1/4}\hat{x}^{1/2},$ (12)
$\epsilon^{n}g_{n}$ $\sim$ $\frac{\epsilon^{1/4}\Lambda(-1)^{n}\Gamma(n-1)}{\hat{x}^{2n-1/2}}.$ (13)
At leading order in (11),
$\hat{g}_{0}-2\hat{x}\hat{g}_{0}^{\prime}+\hat{g}_{0}^{\prime\prime}=0.$
Writing
$\hat{g}_{0}=\sum_{n=0}^{\infty}c_{n}x^{1/2-2n},$
we find
$c_{n}=-\frac{(2n-5/2)(2n-3/2)c_{n-1}}{4n},\qquad c_{0}=1,$
where the latter equation comes from matching with (12). Thus
$c_{n}=-(-1)^{n}\frac{(3/4)_{n-1}(5/4)_{n-1}}{16(2)_{n-1}},$
where $(x)_{n}=\Gamma(x+n)/\Gamma(x)$ is Pochammer’s symbol. Matching with (13) gives
$\Lambda=\lim_{n\rightarrow\infty}\frac{(-1)^{n}c_{n}}{\Gamma(n-1)}=-\frac{1}{16\Gamma(3/4)\Gamma(5/4)}=-\frac{1}{4\sqrt{2}\,\pi}.$
2.4 Boundary layer in the late terms near $X=0$
So far everything we have done has followed the standard approach to finding the late terms of the expansion, as described in [5], for example. In this section we make one crucial observation, which extends this standard approach, and allows us link the two parts of the expansion in (10) and determine $\omega_{n}$.
This observation is that the large-$n$ asymptotic approximation for $g_{n}$ in the outer region is non-uniform, and rearranges when $X$ is small. We can see this directly from the asymptotic behaviour (9), which is singular at $X=0$, while we know that $g_{n}$ is in fact regular at $X=0$.
Thus there is another inner region near the origin, now not in the small-$\epsilon$ expansion of $g$, but in the large-$n$ expansion of $g_{n}$. To examine this inner region we rescale $X$ by setting $X=\xi/n$. Then the equation for $g_{n}$, equation (7), becomes
$\frac{\xi}{n}g_{n}+2\xi\left(1-\frac{\xi}{n}\right)g_{n}^{\prime}=-n\left(\xi g_{n-1}^{\prime}\right)^{\prime}+2\left(1-\frac{\xi}{n}\right)(\omega_{1}g_{n-1}+\cdots+\omega_{n}g_{0}),$ (14)
where is now $\mathrm{d}/\mathrm{d}\xi$. Writing $X=\xi/n$ in (10), the inner limit of the outer is
$g_{n}$ $\sim$ $\frac{\Lambda(-1)^{n}n\Gamma(n-1)}{\xi(1-\xi/n)^{2n-1/2}}+(1-\xi/n)^{1/2}(C_{0}\log n+C_{1}+\log\xi/n)\omega_{n}$ (15)
$\sim$ $\Lambda(-1)^{n}\Gamma(n)\frac{\mathrm{e}^{2\xi}}{\xi}+((C_{0}-1)\log n+C_{1}+\log\xi)\omega_{n}.$
This motivates writing
$\omega_{n}\sim\Omega(-1)^{n}\Gamma(n),\qquad g_{n}\sim H(\xi)\Omega(-1)^{n}\Gamma(n),$
which, on substituting into (14), gives, at leading order,
$(\xi H^{\prime})^{\prime}+2\xi H^{\prime}=2.$
Thus
$H=\alpha_{1}+\alpha_{2}\operatorname{Ei}(2\xi)+\log\xi,$ (16)
where
$\operatorname{Ei}(z)=\int_{-\infty}^{z}\frac{\mathrm{e}^{t}}{t}\,\mathrm{d}t$
is the exponential integral. Now $g_{n}$ should be regular as $\xi\rightarrow 0$. Since $\operatorname{Ei}(2\xi)\sim\log\xi$ as $\xi\rightarrow 0$, we need
$\alpha_{2}=-1$
to remove the logarithmic singularity at $\xi=0$. Since $\operatorname{Ei}(2\xi)$ exhibits Stokes phenomenon for large $\xi$ there will be a switch in the behaviour of the late terms depending on the argument of $\xi$—this is what is known as the higher-order Stokes phenomenon, a Stokes phenomenon not in the asymptotic expansion of $g$ as a function of $\epsilon$, but in the late-term approximation of $g_{n}$ [6, 11]. There is a higher-order Stokes line when $\xi$ crosses the positive real axis, across which the constant contribution to the large-$\xi$ approximation of $H$ (i.e. in the outer limit of the inner expansion) changes. Note that there is no Stokes phenomenon associated with the particular solution $\log\xi$, so that the coefficient of $\mathrm{e}^{2\xi}/\xi$ is fixed. This will not be the case in our other examples.
To complete the analysis and determine $\Omega$ we need to match (16) with (15). As $\xi\rightarrow\infty$
$\operatorname{Ei}(2\xi)\sim\frac{\mathrm{e}^{2\xi}}{2\xi}.$
Matching with (15) gives $C_{0}=1$ and $\Omega=-2\Lambda$ so that
$\omega_{n}\sim-2\Lambda(-1)^{n}\Gamma(n)=\frac{(-1)^{n}\Gamma(n)}{2\sqrt{2}\,\pi},$ (17)
as $n\rightarrow\infty$.
[…]
Figure 1: Divergence of the coefficients in the asymptotic expansion of $\omega$. (a) coefficients determined numerically from (4)-(5). The linear growth is consistent with factorial divergence. (b) The ratio of the numerical value to the asymptotic prediction (17). Blue is the base series, while orange, green and red correspond to enhanced convergence using Richardson extrapolation on two, three and four terms respectively. The black line is the asymptote, included to aid the eye. The convergence is slower than expected because of the presence of log terms in the higher-order corrections, unaccounted for in the extrapolation.
In Fig. 1(b) this result is compared with $\omega_{n}$ found by numerically iterating (4)-(5). The agreement is found to be good, though the convergence is slightly slower than expected because of the presence of log terms in the higher-order corrections.
3 Example 2: Anharmonic oscillator
Having introduced the procedure with a simple model problem, we now consider the classical problem of the anharmonic oscillator [2, 3]. Much of the analysis follows the same framework, though the details of the boundary layer in the late-term approximation analogous to §2.4 are a little different.
Consider [^2] Note the typo in equation (1.1) of [3] in which the minus sign is missing.
[…]
with
$\Psi\rightarrow 0\qquad\mbox{ as }x\rightarrow\pm\infty.$
3.1 Inner region
We first factor out the decay at infinity by writing $\Psi=\mathrm{e}^{-x^{2}/4}g$ to give
$g^{\prime\prime}+xg^{\prime}+\frac{g}{2}+\frac{\epsilon x^{4}g}{4}-\lambda g=0.$ (18)
Now expand
$g=\sum_{n=0}^{\infty}\epsilon^{n}g_{n},\qquad\lambda=\sum_{n=0}^{\infty}\epsilon^{n}\lambda_{n},$ (19)
to give
$-g_{0}^{\prime\prime}+xg_{0}^{\prime}+\frac{g_{0}}{2}-\lambda_{0}g_{0}$ $=$ $0,$ (20)
$-g_{n}^{\prime\prime}+xg_{n}^{\prime}+\frac{g_{n}}{2}-\lambda_{0}g_{n}$ $=$ $-\frac{x^{4}g_{n-1}}{4}+\sum_{k=1}^{n}\lambda_{k}g_{n-k},\qquad n\geq 1.$ (21)
For the first eigenvalue, the leading-order solution is $g_{0}=1$, $\lambda_{0}=1/2$, and, in general
$g_{n}=\sum_{k=1}^{2n}a_{n,k}x^{2k}$ (22)
with
$2ka_{n,k}$ $=$ $(2k+2)(2k+1)a_{n,k+1}-\frac{1}{4}a_{n-1,k-2}+\sum_{i=1}^{n}\lambda_{i}a_{n-i,k},\qquad k=2n,\ldots,1,$ (23)
$-2a_{n,1}$ $=$ $\lambda_{n},$ (24)
with the convention that $a_{n,k}=0$ for $k>2n$ and $k<1$. Equations (23)-(24) are equivalent to eqn. (6.3) in [3]. It is argued in [3] that the leading-order late-term behaviour of (23)-(24) is the same as that of the linearised equation (i.e. with the final sum omitted). With further approximation, and quite a bit of analysis, Bender & Wu manage to extract the leading-order behaviour of $\lambda_{n}$. Here we show how this may be obtained by following the systematic procedure outlined in Section 2.
3.2 Outer region
As before, the expansion (19) is not uniform in xx—it rearranges when $x$ is large. In this section we develop the corresponding expansion valid for large $x$. We subtract off the leading-order eigenvalue by writing
$\lambda=\frac{1}{2}+\epsilon\bar{\lambda}.$ (25)
We rescale into the far field by setting $\epsilon^{1/2}x=X$ to give
$\epsilon^{2}g^{\prime\prime}+\epsilon Xg^{\prime}+\frac{X^{4}g}{4}-\epsilon^{2}\bar{\lambda}g=0.$ (26)
The solution this time is of Liouville-Green (WKB) form,
$g=\mathrm{e}^{\phi/\epsilon}A,\qquad A\sim\sum_{n=0}^{\infty}\epsilon^{n}A_{n}.$ (27)
Substituting (27) into (26) and equating coefficients of powers of $\epsilon$ gives, following a routine calculation,
$\phi=\frac{1}{6}+\frac{X^{2}}{4}-\frac{(1+X^{2})^{3/2}}{6},\qquad A_{0}=\frac{\sqrt{2}}{(1+X^{2})^{1/4}\sqrt{1+\sqrt{1+X^{2}}}},$ (28)
where the normalisation $\sqrt{2}$ comes from matching with the inner region. In general, the equation for $A_{n}$ is
$X(1+X^{2})^{1/2}A_{n}^{\prime}+\left((1+X^{2})^{1/2}-\frac{1}{2}-\frac{1}{2(1+X^{2})^{1/2}}\right)A_{n}-A_{n-1}^{\prime\prime}\\ -\bar{\lambda}_{0}A_{n-1}-\bar{\lambda}_{1}A_{n-2}-\cdots-\bar{\lambda}_{n-1}A_{0}=0.$
As before, there are two sources of divergence in $A_{n}$: the usual factorial/power from repeated differentiation of the singularity in $A_{0}$, and a factorial/constant divergence driven by $\bar{\lambda}_{n}$. For the first, we use the usual factorial/power ansatz following the procedure in [5] to find
$A_{n}\sim\frac{\Lambda}{(1+X^{2})^{1/4}\sqrt{1-\sqrt{1+X^{2}}}}\frac{(-1)^{n}3^{n}\Gamma(n)}{(1+X^{2})^{3n/2}}.$ (29)
The other part of $A_{n}$ satisfies
$X(1+X^{2})^{1/2}A_{n}^{\prime}+\left((1+X^{2})^{1/2}-\frac{1}{2}-\frac{1}{2(1+X^{2})^{1/2}}\right)A_{n}\sim\bar{\lambda}_{n-1}A_{0},$
giving
$A_{n}\sim\frac{\bar{\lambda}_{n-1}}{(1+X^{2})^{1/4}\sqrt{1+\sqrt{1+X^{2}}}}\left(C_{0}\log n+C_{1}-\tanh^{-1}\sqrt{1+X^{2}}\right),$
where, as in §2.2, the presence of a logarithm in $X$ necessitates the inclusion of a logarithm in $n$. Together
$A_{n}\sim\frac{\Lambda}{(1+X^{2})^{1/4}\sqrt{1-\sqrt{1+X^{2}}}}\frac{(-1)^{n}3^{n}\Gamma(n)}{(1+X^{2})^{3n/2}}\\ +\frac{\bar{\lambda}_{n-1}}{(1+X^{2})^{1/4}\sqrt{1+\sqrt{1+X^{2}}}}\left(C_{0}\log n+C_{1}-\tanh^{-1}\sqrt{1+X^{2}}\right).$ (30)
To determine $\Lambda$ we need to match with an inner region in the vicinity of either $X=\mathrm{i}$ or $X=-\mathrm{i}$. This problem is slightly unusual in that there are two singularities in the leading-order solution, but they each produce a late-term behaviour with the same singulant $\chi$, so that there is only one factorial/power divergence in the late terms.
3.3 Inner region near $X=\mathrm{i}$
We set $X=\mathrm{i}-\mathrm{i}\epsilon^{2/3}\hat{x}/2$, $A=\epsilon^{-1/6}\hat{A}$ to give, at leading order,
$2\hat{x}^{1/2}\hat{A}^{\prime}-\frac{\hat{A}}{2\hat{x}^{1/2}}+4\hat{A}^{\prime\prime}=0.$
To match with (28) requires
$\hat{A}\sim\frac{\sqrt{2}}{\hat{x}^{1/4}}\qquad\mbox{ as }\hat{x}\rightarrow\infty$. (31)
Writing
$\hat{A}=\sqrt{2}\sum_{n=0}^{\infty}\frac{g_{n}}{\hat{x}^{1/4+3n/2}},$ (32)
gives
$g_{n}=-\frac{(6n-1)(6n-5)g_{n-1}}{12n},\qquad g_{0}=1,$
where the latter condition comes from (31). Thus
$g_{n}=\frac{5(-1)^{n}3^{n-2}(7/6)_{n-1}(11/6)_{n-1}}{4(2)_{n-1}}.$ (33)
The inner limit of (29) is
$\epsilon^{n}A_{n}\sim\frac{\Lambda}{\epsilon^{1/6}\hat{x}^{1/4}}\frac{(-1)^{n}3^{n}\Gamma(n)}{\hat{x}^{3n/2}}.$ (34)
Matching (34) with (32) gives
$\Lambda=\sqrt{2}\lim_{n\rightarrow\infty}\frac{g_{n}}{(-1)^{n}3^{n}\Gamma(n)}=\frac{5\sqrt{2}}{36\Gamma(7/6)\Gamma(11/6)}=\frac{1}{\pi\sqrt{2}}.$
3.4 Boundary layer in the late terms near $X=0$
Here we come to the key step. The large $n$ asymptotic series for $A_{n}$ in the outer region is nonuniform, and rearranges when $X$ is small—there is another inner region near the origin. We emphasise again that this is a boundary layer not in the small-$\epsilon$ expansion of $A$, but in the large $n$-expansion of $A_{n}$. Again, this nonuniformity is evident because the asymptotic formula (30) is singular at $X=0$, while $A_{n}$ should be regular there. This time the appropriate scaling for the inner region is $X=\xi/n^{1/2}$ (so that $XA_{n}^{\prime}$ balances with $A_{n-1}^{\prime\prime}$), giving
$\xi A_{n}^{\prime}+\frac{3\xi^{2}}{4n}A_{n}-nA_{n-1}^{\prime\prime}+\cdots-\bar{\lambda}_{0}A_{n-1}-\bar{\lambda}_{1}A_{n-2}-\cdots-\bar{\lambda}_{n-1}A_{0}=0.$ (35)
As $X\rightarrow 0$ in (30),
$\epsilon^{n}A_{n}\sim\frac{\sqrt{2}\,\Lambda}{-\mathrm{i}\sqrt{X^{2}}}\frac{(-1)^{n}3^{n}\Gamma(n)}{(1+X^{2})^{3n/2}}+\bar{\lambda}_{n-1}\left(C_{0}\log n+C_{1}\pm\frac{\mathrm{i}\pi}{2}-\log 2+\log X\right).$ (36)
Note that there are two choices of branch to be made here—one for $\sqrt{X^{2}}$ arising from $\sqrt{1-\sqrt{1+X^{2}}}$ and one for the constant $\pm\mathrm{i}\pi/2$ arising from $\tanh^{-1}\sqrt{1+X^{2}}$. In particular note that when matching to find $\Lambda$ we took $\sqrt{1-\sqrt{1+X^{2}}}$ to be real and positive when $X$ approached $\pm\mathrm{i}$, which means we need $\mathrm{i}\sqrt{X^{2}}$ to be real and positive when $X$ is on the imaginary axis; in turn this means we need $\sqrt{X^{2}}=X$ when $X$ is positive imaginary, and $\sqrt{X^{2}}=-X$ when $X$ is negative imaginary. We will return to this choice and the position of the branch cuts shortly, when we match with the inner solution.
With $X=\xi/n^{1/2}$, (36) is
$\epsilon^{n}A_{n}\sim\mathrm{i}\sqrt{2}\,\Lambda n^{1/2}(-1)^{n}3^{n}\Gamma(n)\frac{\mathrm{e}^{-3\xi^{2}/2}}{\sqrt{\xi^{2}}}+\bar{\lambda}_{n-1}\left((C_{0}-1/2)\log n+C_{1}\pm\frac{\mathrm{i}\pi}{2}-\log 2+\log\xi\right).$ (37)
This motivates setting $A_{n}\sim H\Omega(-1)^{n}3^{n}\Gamma(n+1/2)$, $\bar{\lambda}_{n-1}\sim\Omega(-1)^{n}3^{n}\Gamma(n+1/2)$. Using this ansatz in (35) gives
$3\xi H^{\prime}+H^{\prime\prime}-3=0,$
so that
$H=\alpha_{1}+\alpha_{2}\int_{0}^{\xi}\mathrm{e}^{-3u^{2}/2}\,\mathrm{d}u+3\int_{0}^{\xi}\mathrm{e}^{-3t^{2}/2}\int_{0}^{t}\mathrm{e}^{3u^{2}/2}\,\mathrm{d}u\,\mathrm{d}t.$ (38)
Both the particular integral and the complementary function exhibit Stokes phenomenon for large $\xi$, so that there will be a switch in the behaviour of the late terms depending on the argument of $\xi$, corresponding to the higher-order Stokes phenomenon. There is a higher-order Stokes line due to the particular integral when $\xi$ crosses the real axis, across which the coefficient of $\mathrm{e}^{-\xi^{2}}/\xi$ in the far field (i.e. in the outer limit of the inner expansion) changes. The branch cut associated with $\sqrt{\xi^{2}}$ in (37) must be chosen to line up with this higher-order Stokes line. In addition, there is a higher-order Stokes line on the imaginary axis, across which the constant in the far field changes. The branch cut associated with $\pm\mathrm{i}\pi/2$ must be chosen to align with this higher-order Stokes line. As $\xi\rightarrow\infty$ in the first quadrant,
$H\sim\alpha_{1}+\frac{\alpha_{2}\sqrt{\pi}}{\sqrt{6}}+\left(-\frac{\alpha_{2}}{3}-\frac{\mathrm{i}\sqrt{\pi}}{\sqrt{6}}\right)\frac{\mathrm{e}^{-3\xi^{2}/2}}{\xi}+\cdots+\log\xi+\frac{1}{2}(\gamma_{E}+\log 6)+\cdots.$ (39)
where $\gamma_{E}$ is the Euler gamma. As $\xi\rightarrow\infty$ in the third quadrant,
$H\sim\alpha_{1}-\frac{\alpha_{2}\sqrt{\pi}}{\sqrt{6}}+\left(-\frac{\alpha_{2}}{3}+\frac{\mathrm{i}\sqrt{\pi}}{\sqrt{6}}\right)\frac{\mathrm{e}^{-3\xi^{2}/2}}{\xi}+\cdots+\log\xi++\frac{1}{2}(\gamma_{E}+\log 6-\mathrm{i}\pi)+\cdots.$ (40)
In the model problem of Section 2 the key coefficient $\alpha_{2}$ was determined by imposing that $H$ was regular at $\xi=0$. In this case (38) is regular at the origin for all $\alpha_{1}$, $\alpha_{2}$, and it is matching with the outer solution which determines $\alpha_{2}$.
Matching (39) with (37) as $\xi\rightarrow\infty$ in the first quadrant gives
$\left(-\frac{\alpha_{2}}{3}-\frac{\mathrm{i}\sqrt{\pi}}{\sqrt{6}}\right)\Omega=\mathrm{i}\sqrt{2}\Lambda,\qquad C_{0}=1/2,\quad C_{1}+\frac{\mathrm{i}\pi}{2}-\log 2=\alpha_{1}+\frac{\alpha_{2}\sqrt{\pi}}{\sqrt{6}}+\frac{\gamma_{E}}{2}+\log 6.$
Matching (40) with (37) as $\xi\rightarrow\infty$ in the third quadrant gives
$\left(-\frac{\alpha_{2}}{3}+\frac{\mathrm{i}\sqrt{\pi}}{\sqrt{6}}\right)\Omega=-\mathrm{i}\sqrt{2}\Lambda,\qquad C_{0}=1/2,\quad C_{1}-\frac{\mathrm{i}\pi}{2}-\log 2=\alpha_{1}-\frac{\alpha_{2}\sqrt{\pi}}{\sqrt{6}}+\frac{\gamma_{E}}{2}+\log 6-\mathrm{i}\pi.$
Thus $\alpha_{2}=0$ and
$\Omega=-\frac{2\sqrt{3}}{\sqrt{\pi}}\Lambda=-\frac{\sqrt{6}}{\pi^{3/2}}.$
This gives, finally,
$\bar{\lambda}_{n-1}=\lambda_{n}\sim\frac{(-1)^{n+1}\sqrt{6}}{\pi^{3/2}}\,3^{n}\Gamma(n+1/2),$ (41)
in agreement with [3].
[…]
Figure 2: The ratio of the numerical value found by iterating (23)-(24) to the asymptotic prediction (41). Blue is the base series, while orange and green correspond to enhanced convergence using Richardson extrapolation on two and three terms respectively. The black line is the asymptote, included to aid the eye.
In Fig. 2 this result is compared with $\lambda_{n}$ found by numerically iterating (23)-(24); the agreement is excellent.
4 Example 3: Simplified Rossby waves
Our third example is a simplified version of the model problem for an equatorially-trapped Rossby wave considered in [12]. Consider
[…]
with $\psi(0)=1$. Essentially the problem considered in [12] has the $\epsilon^{2}$ in the third term replaced with $\epsilon$. The switch to $\epsilon^{2}$ makes the inner region below more complicated, but significantly simplifies all the other regions of the analysis. This weakens the strength of the pole in Boyd and Natarov’s Hermite-with-pole equation [4], while keeping the same structure.
4.1 Inner region
We will see that the expansion in powers of $\epsilon$ proceeds as
$\psi=\sum_{n=0}^{\infty}\epsilon^{n}\psi_{n},\qquad\lambda=\sum_{n=0}^{\infty}\epsilon^{2n}\lambda_{n}.$ (42)
At leading order
$\psi_{0}^{\prime\prime}-2x\psi_{0}^{\prime}=\lambda_{0}\psi_{0}.$
In order for $\mathrm{e}^{-x^{2}/2}\psi(y)$ to decay as $x\rightarrow\pm\infty$ we need $\psi_{0}$ to be a Hermite polynomial. The leading eigenvalue therefore has $\psi_{0}=1$, $\lambda_{0}=0$. In general
$\psi_{n}^{\prime\prime}-2x\psi_{n}^{\prime}+\sum_{k=2}^{n}(-x)^{k-2}\psi_{n-k}=\sum_{k=0}^{\lfloor n/2\rfloor}\lambda_{k}\psi_{n-2k},\qquad n\geq 1.$
In particular, we find $\psi_{1}=0$, while
$\psi_{2}^{\prime\prime}-2x\psi_{2}^{\prime}+1=\lambda_{1},$
so that $\psi_{2}=0$, $\lambda_{1}=1$. Separating even and odd indices,
$\psi_{2n+1}^{\prime\prime}-2x\psi_{2n+1}^{\prime}+\sum_{k=1}^{n}(-x)^{2k-2}\psi_{2(n-k)+1}+\sum_{k=1}^{n}(-x)^{2k-1}\psi_{2(n-k)}$ $=$ $\sum_{k=0}^{n}\lambda_{k}\psi_{2(n-k)+1},$
$\psi_{2n}^{\prime\prime}-2x\psi_{2n}^{\prime}+\sum_{k=1}^{n}(-x)^{2k-2}\psi_{2(n-k)}+\sum_{k=1}^{n-1}(-x)^{2k-1}\psi_{2(n-k)-1}$ $=$ $\sum_{k=0}^{n}\lambda_{k}\psi_{2(n-k)}.$
The solutions are of the form
$\psi_{2n+1}=\sum_{k=1}^{n}a_{2n+1,k}x^{2k-1},\qquad\psi_{2n}=\sum_{k=0}^{n-1}a_{2n,k}x^{2k},$ (43)
with
$2(2k-1)a_{2n+1,k}$ $=$ $2k(2k+1)a_{2n+1,k+1}+\sum_{m=1}^{n}a_{2(n-m)+1,k-m+1}$ (44)
 $\qquad\mbox{ }-\sum_{m=1}^{n}a_{2(n-m),k-m}-\sum_{m=0}^{n}\lambda_{m}a_{2(n-m)+1,k},$
$4ka_{2n,k}$ $=$ $(2k+2)(2k+1)a_{2n,k+1}+\sum_{m=1}^{n}a_{2(n-m),k-m+1}$ (45)
 $\qquad\mbox{ }-\sum_{m=1}^{n-1}a_{2(n-m)-1,k-m+1}-\sum_{m=0}^{n}\lambda_{m}a_{2(n-m),k},$
with $a_{n,0}=0$ for $n>0$ and the convention that $a_{2n+1,k}=0$ and $a_{2n,k}=0$ if $k>n$ or $k<1$. For each $n$ equations (44)-(45) may be solved iteratively stepping down from $k=n$. The solvability condition determining $\lambda_{n}$ comes from setting $k=0$ in (45), giving $\lambda_{n}=2a_{2n,1}$. We now follow the procedure of §2 to determine the asymptotic behaviour of $\lambda_{n}$ as $n\rightarrow\infty$.
4.2 Outer region
As usual, the expansion (42) is not uniform in $x$ and rearranges when $x$ is large. We set $\epsilon x=X$ to give
[…]
The outer expansion now proceeds straightforwardly in powers of $\epsilon^{2}$ as
$\psi=\sum_{n=0}^{\infty}\epsilon^{2n}\psi_{n},\qquad\lambda=\sum_{n=0}^{\infty}\epsilon^{2n}\lambda_{n}.$ (46)
At leading order
[…]
with solution
$\psi_{0}=B_{0}X^{-\lambda_{0}/2}.$
For there to be no singularity at $X=0$ requires $\lambda_{0}/2$ to be a non-positive integer, in agreement with the inner analysis in §4.1. The leading eigenvalue therefore has $\lambda_{0}=0$, $\psi_{0}=1$. At next order
[…]
with solution
$\psi_{1}=B_{1}+\frac{(1-\lambda_{1})}{2}\log X-\frac{1}{2}\log(1+X).$
For there to be no singularity at $X=0$ we require $\lambda_{1}=1$, in agreement with §4.1. The boundary condition $\psi_{1}(0)=0$ (more properly a matching condition with the inner region) gives $B_{1}=0$. In general
[…]
As usual, there are two types of divergence: a factorial/power from repeated differentiation of the singularity $\log(1+X)$ in $\psi_{1}$, and a factorial/constant from $\lambda_{n}$. For the first, we use the usual ansatz to give
$\psi_{n}\sim\frac{\Lambda\Gamma(n-1)}{X(1-X^{2})^{n-1}}.$ (48)
Here we see a curious feature of this example—the late term behaviour $\psi_{n}$ was driven by a singularity at X=1X=-1, but the singulant vanishes also at $X=+1$. Whereas in the anharmonic oscillator problem of §3 both singularities $X=\pm\mathrm{i}$ were present in the early terms, here only X=1X=-1 is present in the early terms. The resolution of this apparent paradox, as described in [12], is a higher-order Stokes line which turns off the contribution (48) in a region enclosing $X=1$. We will return to this point later when matching with an inner region near $X=0$.
The other part of $\psi_{n}$ satisfies $\psi_{n}\sim Q\lambda_{n}$ where
$2XQ^{\prime}=1,$
giving
$Q=C-\frac{1}{2}\log X.$
As usual, the presence of a logarithm means we need to modify the large $n$ ansatz to essentially allow $C$ to depend on $\log n$. Putting both parts of $\psi_{n}$ together we have
$\psi_{n}\sim\frac{\Lambda\Gamma(n-1)}{X(1-X^{2})^{n-1}}+\left(-\frac{1}{2}\log X+C_{0}\log n+C_{1}\right)\lambda_{n}.$ (49)
The next step is to determine the constant $\Lambda$, by matching with an inner region near the singularity at X=1X=-1.
4.3 Inner region near X=1X=-1
We set $X=-1+\epsilon^{2}\hat{x}$. Then the inner limit of the outer expansion satisfies
$\psi_{0}+\epsilon^{2}\psi_{1}$ $\sim$ $1-\epsilon^{2}\log\epsilon-\frac{\epsilon^{2}}{2}\log\hat{x}$, (50)
$\epsilon^{2n}\psi_{n}$ $\sim$ $-\frac{\epsilon^{2}\Lambda\Gamma(n-1)}{(2\hat{x})^{n-1}}.$ (51)
Equation (50) motivates writing $\psi=1-\epsilon^{2}\log\epsilon+\epsilon^{2}\hat{\psi}$ to give the inner equation as
[…]
At leading order
[…]
Thus
$\hat{\psi}_{0}=\beta_{1}+\beta_{2}\mathrm{e}^{-2\hat{x}}-\frac{1}{2}\log\hat{x}+\frac{1}{2}\mathrm{e}^{-2\hat{x}}\operatorname{Ei}(2\hat{x})=\beta_{1}+\beta_{2}\mathrm{e}^{-2\hat{x}}-\frac{1}{2}\log\hat{x}+\frac{1}{2}\sum_{n=0}^{\infty}\frac{\Gamma(n+1)}{(2\hat{x})^{n+1}}.$
Matching with (51) gives
$\Lambda=-\frac{1}{2}.$
4.4 Boundary layer in the late terms near $X=0$
[…]
The inner limit of the outer expansion (49) is, for $X<0$,
$\psi_{n}$ $\sim$ $-\frac{\Gamma(n-1)n^{1/2}}{2\xi(1-\xi^{2}/n)^{n-1}}+\left(-\frac{1}{2}\log(\xi)+C\right)\lambda_{n}$ (53)
$\sim$ $-\frac{\Gamma(n-1/2)}{2}\frac{\mathrm{e}^{\xi^{2}}}{\xi}+\left(-\frac{1}{2}\log(\xi)+C_{0}\log n+C_{1}\right)\lambda_{n}.$
This motivates setting $\lambda_{n}=\Omega\Gamma(n-1/2)$, $\psi_{n}\sim H\Omega\Gamma(n-1/2)$ in (52), to give
$H^{\prime\prime}-2\xi H^{\prime}=1,$
[…]
$H=\alpha_{1}+\alpha_{2}\int_{0}^{\xi}\mathrm{e}^{t^{2}}\,\mathrm{d}t+\int_{0}^{\xi}\mathrm{e}^{t^{2}}\int_{0}^{t}\mathrm{e}^{-p^{2}}\,\mathrm{d}p\,\mathrm{d}t.$
Both the particular integral and the complementary function exhibit Stokes phenomenon for large $\xi$, corresponding to the higher-order Stokes phenomenon in $\psi$. There is a higher-order Stokes line due to the particular integral when $\xi$ crosses the imaginary axis, across which the coefficient of $\mathrm{e}^{\xi^{2}}/\xi$ in the far field changes. In addition, there is a higher-order Stokes line on the real axis, across which the constant in the far field changes. For real $\xi$, as $\xi\rightarrow$-$\infty$
$H\sim\alpha_{1}+\mathrm{i}\frac{\alpha_{2}\sqrt{\pi}}{2}+\left(\frac{\alpha_{2}}{2}-\frac{\sqrt{\pi}}{4}\right)\frac{\mathrm{e}^{\xi^{2}}}{\xi}+\cdots-\frac{1}{4}\left(\gamma_{E}+\log(-4\xi^{2})+\cdots\right)$
while as $\xi\rightarrow+\infty$
$H\sim\alpha_{1}-\mathrm{i}\frac{\alpha_{2}\sqrt{\pi}}{2}+\left(\frac{\alpha_{2}}{2}+\frac{\sqrt{\pi}}{4}\right)\frac{\mathrm{e}^{\xi^{2}}}{\xi}+\cdots-\frac{1}{4}\left(\gamma_{E}+\log(-4\xi^{2})+\cdots\right),$
where $\gamma_{E}$ is the Euler gamma.
4.5 Matching with the outer
The outer limit of the inner is
$\psi_{n}$ $\sim$ $\Omega\left(\frac{\alpha_{2}}{2}-\frac{\sqrt{\pi}}{4}\right)\frac{\mathrm{e}^{\xi^{2}}}{\xi}\Gamma(n-1/2)+\cdots-\frac{\lambda_{n}}{2}\left(\log\xi+\cdots\right)\qquad\mbox{ as }\xi\rightarrow-\infty,$
$\psi_{n}$ $\sim$ $\Omega\left(\frac{\alpha_{2}}{2}+\frac{\sqrt{\pi}}{4}\right)\frac{\mathrm{e}^{\xi^{2}}}{\xi}\Gamma(n-1/2)+\cdots-\frac{\lambda_{n}}{2}\left(\log\xi+\cdots\right)\qquad\mbox{ as }\xi\rightarrow+\infty.$
Matching with (53) as $\xi\rightarrow$-$\infty$ gives
$\Omega\left(\frac{\alpha_{2}}{2}-\frac{\sqrt{\pi}}{4}\right)=-\frac{1}{2}.$
For $X>0$, as per the discussion following (48), there can be no exponential term in the outer, because there must be no singularity at $X=1$. Thus, matching as $\xi\rightarrow\infty$ gives
$\Omega\left(\frac{\alpha_{2}}{2}+\frac{\sqrt{\pi}}{4}\right)=0.$
Together
$\alpha_{2}=-\frac{\sqrt{\pi}}{2},\qquad\Omega=\frac{1}{\sqrt{\pi}},$
so that
$\lambda_{n}\sim\frac{\Gamma(n-1/2)}{\sqrt{\pi}}.$ (54)
In Fig. 3 this result is compared with $\lambda_{n}$ found by numerically iterating (44)-(45); the agreement is good, though the convergence is slower than expected because of the presence of logarithmic terms in the higher-order corrections.
[…]
Figure 3: The ratio of the numerical value found by iterating (44)-(45) to the asymptotic prediction (54). Blue is the base series, while orange, green and red correspond to enhanced convergence using Richardson extrapolation on two, three and four terms respectively. The black line is the asymptote, included to aid the eye. The convergence is slower than expected because of the presence of logarithmic terms in the higher-order corrections, unaccounted for in the extrapolation.
5 Example 4: Divergence driven by two singularities
Our final example is chosen to illustrate that the divergence of the eigenvalue can be driven by more than one singularity in the outer solution, leading to more exotic behaviour. This is exactly what happens in the model in [7] concerning quasinormal modes of Reissner-Nordström-de Sitter black holes. The form of this divergence is more difficult to pick up with other methods, and the interaction between two singularities makes it difficult to guess the form of the divergence from numerical calculations of the leading terms in the series.
Consider, as a model problem,
$\frac{b^{2}+(c+\epsilon x)^{2}}{b^{2}+c(c+\epsilon x)}\left(-\omega g+xg^{\prime}\right)+g+(xg^{\prime})^{\prime}=0,\qquad-\infty<x<0,$ (55)
with
$g(0)=1,\qquad g(x)=o(\mathrm{e}^{-x})\mbox{ as }x\rightarrow\infty.$
The relationship with (1) is clear—the coefficient of the first term has been modified to generate an outer solution with two singularities.
5.1 Inner region
We expand
$g=\sum_{n=0}^{\infty}\epsilon^{n}g_{n},\qquad\omega=\sum_{n=0}^{\infty}\epsilon^{n}\omega_{n},$
to give at leading order
$\omega_{0}g_{0}+xg_{0}^{\prime}+g_{0}+(xg_{0}^{\prime})^{\prime}=0,$
with solution $g_{0}=1$, $\omega_{0}=1$. At next order
$\frac{cx}{b^{2}+c^{2}}+\left(-\omega_{1}+xg_{1}^{\prime}\right)+(xg_{1}^{\prime})^{\prime}=0,$
with solution
$g_{1}=-\frac{cx}{b^{2}+c^{2}},\qquad\omega_{1}=\frac{c}{b^{2}+c^{2}}.$
In general
$g_{n}=\sum_{i=1}^{n}a_{ni}x^{i}.$
with
$ia_{ni}=\sum_{k=1}^{n}\omega_{k}a_{n-k,i}-(i+1)^{2}a_{n,i+1}\\ -\frac{c}{(b^{2}+c^{2})}\left(-2\sum_{k=0}^{n-1}\omega_{k}a_{n-k-1,i-1}+2(i-1)a_{n-1,i-1}+a_{n-1,i-1}+i^{2}a_{n-1,i}\right)\\ -\frac{1}{(b^{2}+c^{2})}\left(-\sum_{k=0}^{n-2}\omega_{k}a_{n-2-k,i-2}+(i-2)a_{n-2,i-2}\right),$ (56)
and $\omega_{n}=a_{n,1}$. As usual, we can iterate (56) numerically. Figure 4 shows $|\omega_{n}|^{1/n}$ as a function of $n$; the linear growth in $n$ is consistent with factorial growth in $\omega_{n}$ at large $n$. Note that this growth is not nearly as smooth as that in Fig. 1(a), with some ripples present. Similar ripples can be seen in Fig. 2 of [7]. These ripples are a direct result of the interaction of the contributions from the two singularities in the outer problem.
[…]
(a) $b=1$, $c=1$
[…]
(b) $b=1$, $c=3$
Figure 4: Divergence of the coefficients in the asymptotic expansion of $\omega$, determined numerically from (56). The linear growth is consistent with factorial divergence.
5.2 Outer region
We set $\epsilon x=X$ to give
$\frac{b^{2}+(c+X)^{2}}{b^{2}+c(c+X)}\left(-\omega g+Xg^{\prime}\right)+g+\epsilon(Xg^{\prime})^{\prime}=0.$
Expanding
$g=\sum_{n=0}^{\infty}\epsilon^{n}g_{n},\qquad\omega=\sum_{n=0}^{\infty}\epsilon^{n}\omega_{n},$ (57)
and using $\omega_{0}=1$ gives
$\frac{b^{2}+(c+X)^{2}}{b^{2}+c(c+X)}\left(-g_{0}+Xg_{0}^{\prime}\right)+g_{0}=0,$
with solution
$g_{0}=\frac{\sqrt{b^{2}+(c+X)^{2}}}{\sqrt{b^{2}+c^{2}}},$ (58)
where we have used the fact that $g_{0}\rightarrow 1$ as $X\rightarrow 0$. We see that $g_{0}$ has singularities when $b^{2}+(c+X)^{2}=0$, i.e. $X=-c\pm\mathrm{i}b$ (of course, the coefficient of the first term in (55) was chosen to make this the case). In general
$\frac{b^{2}+(c+X)^{2}}{b^{2}+c(c+X)}(-g_{n}+Xg_{n}^{\prime})+g_{n}=-(Xg_{n-1}^{\prime})^{\prime}+\frac{b^{2}+(c+X)^{2}}{b^{2}+c(c+X)}(\omega_{1}g_{n-1}+\cdots+\omega_{n}g_{0}).$
As usual there are two types of divergence: a factorial/power from the differentiation, and a factorial/constant from the $\omega_{n}$. For the first, we use the usual ansatz
$g_{n}=\frac{G\Gamma(n+\gamma)}{\chi^{n+\gamma}}.$ (59)
At leading order in $n$ this gives
$\frac{b^{2}+(c+X)^{2}}{b^{2}+c(c+X)}=-\chi^{\prime}$
so that
$\chi=\frac{(c+X)(cX+c^{2}-2b^{2})}{2c^{2}}+\frac{b^{2}(b^{2}+c^{2})\log(b^{2}+c^{2}+cX)}{c^{3}}+\mbox{const.}$
This time there are two possible late term divergences, one corresponding to each of the two singularities of the leading order solution:
$\chi_{1}$ $=$ $\frac{2\mathrm{i}b^{3}+c(c+X)^{2}-b^{2}(c+2X)}{2c^{2}}+\frac{b^{2}(b^{2}+c^{2})}{c^{3}}\log\left(\frac{b^{2}+c(c+X)}{b^{2}+\mathrm{i}bc}\right),$
$\chi_{2}$ $=$ $\frac{-2\mathrm{i}b^{3}+c(c+X)^{2}-b^{2}(c+2X)}{2c^{2}}+\frac{b^{2}(b^{2}+c^{2})}{c^{3}}\log\left(\frac{b^{2}+c(c+X)}{b^{2}-\mathrm{i}bc}\right).$
At next order,
$\frac{b^{2}+(c+X)^{2}}{b^{2}+c(c+X)}(-G+XG^{\prime})+G$ $=$ $G\chi^{\prime}+X(2G^{\prime}\chi^{\prime}+G\chi^{\prime\prime}).$
Thus
$\frac{(b^{4}+c(c+X)^{2}(c+3X)+b^{2}(2c^{2}+5cX+4X^{2}))}{(b^{2}+c(c+X))^{2}}G+\frac{X(b^{2}+(c+X)^{2})}{b^{2}+c(c+X)}G^{\prime}=0,$
giving
$G=\frac{\Lambda(b^{2}+c(c+X))}{X(b^{2}+(c+X)^{2})^{3/2}}.$
Thus this part of $g_{n}$ satisfies
$g_{n}\sim\frac{\Lambda_{1}(b^{2}+c(c+X))}{X(b^{2}+(c+X)^{2})^{3/2}}\frac{\Gamma(n+\gamma_{1})}{\chi_{1}^{n+\gamma_{1}}}+\frac{\Lambda_{2}(b^{2}+c(c+X))}{X(b^{2}+(c+X)^{2})^{3/2}}\frac{\Gamma(n+\gamma_{2})}{\chi_{2}^{n+\gamma_{2}}}.$
To determine $\gamma_{1}$ and $\gamma_{2}$ we match the order of the singularity as $X\rightarrow-c\pm\mathrm{i}b$ with the early terms. As $X\rightarrow-c+\mathrm{i}b$,
$\chi_{1}\sim\frac{(X+c-\mathrm{i}b)^{2}}{c-\mathrm{i}b},$
$g_{n}\sim-\frac{\mathrm{i}b}{2\sqrt{2}(\mathrm{i}b)^{3/2}(X+c-\mathrm{i}b)^{3/2}}\frac{(c-\mathrm{i}b)^{n+\gamma_{1}}}{(X+c-\mathrm{i}b)^{2n+2\gamma_{1}}}\Lambda_{1}\Gamma(n+\gamma_{1}).$
Comparing powers of $X+c-\mathrm{i}b$ with $g_{0}$ gives
$\frac{3}{2}-2\gamma_{1}=\frac{1}{2}\qquad\Rightarrow\qquad\gamma_{1}=-1.$
A similar comparison as $X\rightarrow-c-\mathrm{i}b$ gives $\gamma_{2}=-1$ also, so that
$g_{n}\sim\frac{\Lambda_{1}(b^{2}+c(c+X))}{X(b^{2}+(c+X)^{2})^{3/2}}\frac{\Gamma(n-1)}{\chi_{1}^{n-1}}+\frac{\Lambda_{2}(b^{2}+c(c+X))}{X(b^{2}+(c+X)^{2})^{3/2}}\frac{\Gamma(n-1)}{\chi_{2}^{n-1}}.$
The other part of $g_{n}$ satisfies $g_{n}=(Q_{0}\log n+Q_{1})\omega_{n}$ where
$\frac{b^{2}+(c+X)^{2}}{b^{2}+c(c+X)}(-Q_{0}+XQ_{0}^{\prime})+Q_{0}$ $=$ $0,$
$\frac{b^{2}+(c+X)^{2}}{b^{2}+c(c+X)}(-Q_{1}+XQ_{1}^{\prime})+Q_{1}$ $=$ $\frac{b^{2}+(c+X)^{2}}{b^{2}+c(c+X)}g_{0},$
giving
$Q_{0}\log n+Q_{1}=\frac{\sqrt{b^{2}+(c+X)^{2}}}{b^{2}+c^{2}}\left(\log X+C_{0}\log n+C_{1}\right).$
Together
$g_{n}\sim\frac{\Lambda_{1}(b^{2}+c(c+X))}{X(b^{2}+(c+X)^{2})^{3/2}}\frac{\Gamma(n-1)}{\chi_{1}^{n-1}}+\frac{\Lambda_{2}(b^{2}+c(c+X))}{X(b^{2}+(c+X)^{2})^{3/2}}\frac{\Gamma(n-1)}{\chi_{2}^{n-1}}\\ +\frac{\sqrt{b^{2}+(c+X)^{2}}}{b^{2}+c^{2}}\left(\log X+C_{0}\log n+C_{1}\right)\omega_{n}.$ (60)
The next step is to determine $\Lambda_{1}$ and $\Lambda_{2}$ through matching with inner regions near $X=-c\pm\mathrm{i}b$.
5.3 Inner region near $X=-c+\mathrm{i}b$
To determine $\Lambda_{1}$ we look near $X=-c+\mathrm{i}b$. We set $X=-c+\mathrm{i}b+\epsilon^{1/2}(c-\mathrm{i}b)^{1/2}\hat{x}$, $g=\epsilon^{1/4}\hat{g}$ to give
$\frac{2\mathrm{i}\epsilon^{1/2}(c-\mathrm{i}b)^{1/2}\hat{x}}{b+\mathrm{i}c}\left(-\omega\hat{g}+\frac{1}{(c-\mathrm{i}b)^{1/2}\epsilon^{1/2}}(-c+\mathrm{i}b+\epsilon^{1/2}(c-\mathrm{i}b)^{1/2}\hat{x})\hat{g}^{\prime}\right)\\ +\hat{g}+\epsilon\left(\frac{1}{(c-\mathrm{i}b)^{1/2}\epsilon^{1/2}}\hat{g}^{\prime}+(-c+\mathrm{i}b+(c-\mathrm{i}b)^{1/2}\epsilon^{1/2}\hat{x})\frac{1}{(c-\mathrm{i}b)\epsilon}\hat{g}^{\prime\prime}\right)=0.$
At leading order
$2\hat{x}\hat{g}_{0}^{\prime}+\hat{g}_{0}-\hat{g}_{0}^{\prime\prime}=0.$
Writing
$\hat{g}_{0}=\frac{\sqrt{2}(\mathrm{i}b)^{1/2}(c-\mathrm{i}b)^{1/4}}{\sqrt{b^{2}+c^{2}}}\sum_{n=0}^{\infty}c_{n}x^{1/2-2n},$ (61)
gives
$c_{n}=\frac{(2n-5/2)(2n-3/2)c_{n-1}}{4n},\qquad c_{0}=1,$
where the latter condition comes from matching with (58). Thus
$c_{n}=-\frac{(3/4)_{n-1}(5/4)_{n-1}}{16(2)_{n-1}}.$
The inner limit of the outer expansion is
$\epsilon^{n}g_{n}$ $\sim$ $-\frac{\mathrm{i}b\epsilon^{n}}{2\sqrt{2}(\mathrm{i}b)^{3/2}(X+c-\mathrm{i}b)^{3/2}}\frac{(c-\mathrm{i}b)^{n-1}}{(X+c-\mathrm{i}b)^{2n-2}}\Lambda_{1}\Gamma(n-1)$ (62)
$\sim$ $-\frac{\epsilon^{1/4}}{2\sqrt{2}(\mathrm{i}b)^{1/2}}\frac{\Lambda_{1}\Gamma(n-1)\hat{x}^{1/2-2n}}{(c-\mathrm{i}b)^{3/4}}.$
Matching (61) with (62) gives
$\Lambda_{1}$ $=$ $-\frac{\sqrt{2}(\mathrm{i}b)^{1/2}(c-\mathrm{i}b)^{1/4}}{\sqrt{b^{2}+c^{2}}}2\sqrt{2}(\mathrm{i}b)^{1/2}(c-\mathrm{i}b)^{3/4}\lim_{n\rightarrow\infty}\frac{c_{n}}{\Gamma(n-1)}$
$=$ $\frac{(\mathrm{i}b)(c-\mathrm{i}b)^{1/2}}{(c+\mathrm{i}b)^{1/2}}\frac{4}{16\Gamma(3/4)\Gamma(5/4)}=\frac{(\mathrm{i}b)(c-\mathrm{i}b)^{1/2}}{\sqrt{2}\,\pi(c+\mathrm{i}b)^{1/2}}.$
A similar calculation near $X=-c-\mathrm{i}b$ shows
$\Lambda_{2}=\frac{(-\mathrm{i}b)(c+\mathrm{i}b)^{1/2}}{\sqrt{2}\,\pi(c-\mathrm{i}b)^{1/2}}=\bar{\Lambda}_{1},$
where an overbar denotes complex conjugation.
5.4 Boundary layer in the late terms near $X=0$
As in all our examples, the large $n$ asymptotic series for $g_{n}$ in the outer region rearranges when $X$ is small, so that there is an inner region near the origin. As in §2 the appropriate rescaling is $X=\xi/n$, under which the equation for $g_{n}$ becomes
$\frac{b^{2}+(c+\xi/n)^{2}}{b^{2}+c(c+\xi/n)}(-g_{n}+\xi g_{n}^{\prime})+g_{n}=-n(\xi g_{n-1}^{\prime})^{\prime}+\frac{b^{2}+(c+\xi/n)^{2}}{b^{2}+c(c+\xi/n)}(\omega_{1}g_{n-1}+\cdots+\omega_{n}g_{0}).$ (63)
Writing $\chi_{1}$ and $\chi_{2}$ in terms of $\xi$ gives
$\chi_{1}$ $\sim$ $\frac{2\mathrm{i}b^{3}-b^{2}c+c^{3}}{2c^{2}}+\frac{b^{2}(b^{2}+c^{2})}{c^{3}}\log\left(1-\frac{\mathrm{i}c}{b}\right)+\frac{\xi}{n}+\cdots=\chi_{0}+\frac{\xi}{n},$
$\chi_{2}$ $\sim$ $\frac{-2\mathrm{i}b^{3}-b^{2}c+c^{3}}{2c^{2}}+\frac{b^{2}(b^{2}+c^{2})}{c^{3}}\log\left(1+\frac{\mathrm{i}c}{b}\right)+\frac{\xi}{n}+\cdots=\bar{\chi}_{0}+\frac{\xi}{n},$
say. Thus the inner limit of the outer solution is
$g_{n}$ $\sim$ $\frac{\Lambda_{1}}{\xi(b^{2}+c^{2})^{1/2}}\frac{\Gamma(n)}{(\chi_{0}+\xi/n)^{n-1}}+\frac{\bar{\Lambda}_{1}}{\xi(b^{2}+c^{2})^{1/2}}\frac{\Gamma(n)}{(\bar{\chi_{0}}+\xi/n)^{n-1}}$ (64)
 $\hskip 170.71652pt\mbox{ }+\left(\log\xi/n+C_{0}\log n+C_{1}\right)\omega_{n}$
$\sim$ $\frac{\Lambda_{1}}{\xi(b^{2}+c^{2})^{1/2}}\frac{\Gamma(n)}{\chi_{0}^{n-1}}\mathrm{e}^{-\xi/\chi_{0}}+\frac{\bar{\Lambda}_{1}}{\xi(b^{2}+c^{2})^{1/2}}\frac{\Gamma(n)}{\bar{\chi_{0}}^{n-1}}\mathrm{e}^{-\xi/\bar{\chi}_{0}}$
 $\hskip 170.71652pt\mbox{ }+\left(\log\xi+(C_{0}-1)\log n+C_{1}\right)\omega_{n}$
From our analyses in §24 we have seen that the boundary-layer approximation to $g_{n}$ comprises a particular integral driven by $\omega_{n}$ and a complementary function matching with the remaining factorial/power divergence of the outer expansion. For the current problem we write the particular integral as $g_{n}=H\omega_{n}$, giving
$(\xi H^{\prime})^{\prime}+\xi H^{\prime}=1,$
so that
$H=\log\xi$
as in §2. Matching this particular solution with (64) gives $C_{0}=1$, $C_{1}=0$. The homogeneous solution may be written
$g_{n}=G(\xi)\frac{\Gamma(n)}{\chi_{0}^{n}}+\bar{G}(\xi)\frac{\Gamma(n)}{\bar{\chi}_{0}^{n}},$
where
$\chi_{0}(\xi G^{\prime})^{\prime}+\xi G^{\prime}=0,$
giving
$G=\alpha_{1}+\alpha_{2}\operatorname{Ei}(-\xi/\chi_{0}).$
Thus together we have
$g_{n}\sim\left(\alpha_{1}+\alpha_{2}\operatorname{Ei}(-\xi/\chi_{0})\right)\frac{\Gamma(n)}{\chi_{0}^{n}}+\left(\bar{\alpha}_{1}+\bar{\alpha}_{2}\operatorname{Ei}(-\xi/\bar{\chi}_{0})\right)\frac{\Gamma(n)}{\bar{\chi}_{0}^{n}}+\omega_{n}\log\xi.$
Now, as in §2, $g_{n}$ should be regular as $\xi\rightarrow 0$. Thus, since $\operatorname{Ei}(\xi)\sim\log\xi$ as $\xi\rightarrow 0$, we need
$\omega_{n}\sim-\alpha_{2}\frac{\Gamma(n)}{\chi_{0}^{n}}-\bar{\alpha}_{2}\frac{\Gamma(n)}{\bar{\chi}_{0}^{n}}.$
To complete the analysis we need to match with (64) to determine $\alpha_{2}$. As $\xi\rightarrow\infty$
$\operatorname{Ei}(-\xi/\chi_{0})\sim-\frac{\chi_{0}\mathrm{e}^{-\xi/\chi_{0}}}{\xi}.$
Thus the outer limit of the inner is
$g_{n}\sim\left(\alpha_{1}-\alpha_{2}\frac{\chi_{0}\mathrm{e}^{-\xi/\chi_{0}}}{\xi}\right)\frac{\Gamma(n)}{\chi_{0}^{n-1}}+\left(\bar{\alpha}_{1}-\bar{\alpha}_{2}\frac{\bar{\chi}_{0}\mathrm{e}^{-\xi/\bar{\chi}_{0}}}{\xi})\right)\frac{\Gamma(n)}{\bar{\chi}_{0}^{n-1}}+\omega_{n}\log\xi.$
Matching with (64) gives
$\alpha_{2}=-\frac{\Lambda_{1}}{(b^{2}+c^{2})^{1/2}}=-\frac{\mathrm{i}b}{\sqrt{2}\,\pi(c+\mathrm{i}b)}.$
Thus
$\omega_{n}\sim\frac{\mathrm{i}b}{\sqrt{2}\,\pi(c+\mathrm{i}b)}\frac{\Gamma(n)}{\chi_{0}^{n}}-\frac{\mathrm{i}b}{\sqrt{2}\,\pi(c-\mathrm{i}b)}\frac{\Gamma(n)}{\bar{\chi}_{0}^{n}}.$ (65)
In Fig. 5 this result is compared with $\omega_{n}$ found by numerically iterating (56) for various values of $b$ and $c$. The sinusoidal oscillation predicted by (65) is clear in Figs. 5(b) and 5(d), when the period of the oscillation is long enough that there are many integers per cycle, but when the period is short $\omega_{n}$ seems to jump around between different longwave oscillations because of aliasing.
[…]
(a) $b=3$, $c=1$
[…]
(b) $b=1$, $c=3$
[…]
(c) $b=1$, $c=1$
[…]
(d) $b=1$, $c=2$
Figure 5: A comparison of the asymptotic approximation (65) with $\omega_{n}$ found by numerically iterating (56), for various values of $b$ and $c$. We normalise by $\Gamma(n)/|\chi_{0}|^{n}$ to remove the exponential growth. The solid curve shows (65) as a continuous function of $n$, which is a sinusoidal oscillation of period $\arg(\chi_{0})/2\pi$. The green dots show (65) evaluated at integer $n$. The red dots are the numerical values.
6 Conclusion
Through four examples, we have demonstrated a systematic procedure for calculating the precise asymptotic behaviour of the late terms of the asymptotic expansion of the eigenvalue in a variety of linear eigenvalue problems. The framework in each of our examples is the same.
After a regular perturbation expansion the eigenfunction at each order is a polynomial, leading to a set of recurrence relations for the coefficients of these polynomials and the coefficients of the eigenvalue expansion. While these relations are easy to iterate numerically to get the leading terms of the eigenvalue expansion, it is hard to extract the late term behaviour from them.
This regular perturbation expansion is nonuniform, and rearranges when $x$ is large. Rescaling to an outer variable the corresponding outer solution can be found, again as an asymptotic power series. This series has a standard factorial/power divergence driven by singularities in the leading-order approximation, and an additional divergence driven by the divergent eigenvalue expansion. In contrast to the original expansion, the late terms in this outer asymptotic expansion are easy to find, using the usual factorial/power ansatz, but the divergence of the eigenvalue is still undetermined.
However, the late term approximation of this outer expansion also non-uniform, now not as $\epsilon\rightarrow 0$ but as $n\rightarrow\infty$. By introducing a local variable in the equation for the late terms of the outer expansion, a new inner expansion is generated in which the two parts of the divergence become coupled, and the eigenvalue is determined.
We hope our framework provides a template by which similar problems of interest may be solved.
Acknowledgements
I would like to acknowledge the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme ‘Applicable resurgent asymptotics: towards a universal theory’, where part of the work on Examples 1 and 3 was undertaken, as well as useful discussions with many of the participants. I would like to single out in particular Ines Aniceto, who among other things introduced me to reference [7] during this programme. This programme was supported by EPSRC grant no. EP/R014604/1.
I would also like to acknowledge the hospitality of the University of Sydney and the support of Australian Research Council Discovery Project DP240101666.
For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.